# Long-time soliton dynamics via a coarse-grained space-time method

### v2

@@ IV Discretization of 1+1D field theory with DEC-QED
In this section, we discuss a discretization procedure for the 1+1D field equations to enable their symplectic integration. Although in this section we will focus on the derivation of the coarse-grained version of the perturbed sine-Gordon equation in space-time, the process for coarse-graining other nonlinear field equations follows a similar procedure (see e.g. Ref. [43]). We consider a primal rectangular grid $M$ that discretizes the space-time computational domain, in which $\Delta x$ and $\Delta t$ are spacings between the neighboring grid points along $x$ and $t$, respectively. We label the vertices, edges, and elemental faces of this mesh as $v$, $e$ and $f$, respectively. Additionally, consider a dual mesh whose vertices are the circumcenters of the rectangular cells $f$ of the primal grid. Each edge of this dual mesh is constructed by connecting the two dual nodes corresponding to two neighboring primal faces (faces that share a primal edge). The dual faces, edges, and nodes are labeled $v^{\dagger}$, $e^{\dagger}$, and $f^{\dagger}$, respectively. Fig. 2a presents a schematic of the spacetime grid, where the primal mesh is colored orange and the dual mesh is in green.
By introducing the 1+1D Minkowski metric tensor $g_{\mu\nu}$ [60, 61], Eq. 2 can be conveniently rewritten as
@@
[…]
@@
where $\bar{\phi_{t}}$ is the average of the field values on the two $\phi_{t}$ edges connected to $v$. Given initial conditions on either $\phi_{t}$ and $\varphi$ everywhere in space or $\phi_{t}$ and $\phi_{x}$ everywhere combined with the value for $\phi$ at the boundary, Eq. 13 can be used to propagate the system in time to study the evolution of $\phi_{t}$, $\phi_{x}$. The procedure is as follows: by labeling each primal vertex using their discrete spatial and temporal indices $i$ and $j$ and each primal edge by the two vertices connected to it, Eq. 13 when applied to a vertex $v(i,j)$ can be rewritten as
@@
$\varphi(i,j\!+\!1)=\phi_{t}[(i,j),(i,j\!+\!1)]\!+\!\varphi(i,j),$ (15)
$\phi_{x}[(i,j\!+\!1),\!(i\!+\!1,j\!+\!1)]=\varphi(i\!+\!1,j\!+\!1)\!-\!\varphi(i,j\!+\!1).$ (16)
Eqs. IV, 15, and 16, along with initial and boundary conditions allow us to determine the fields $\phi_{t},\phi_{x},$ and $\varphi$ everywhere on the spacetime grid. Note that Eq. 13 is effectively a first-order equation in terms of the edge fields $\phi_{x}$ and $\phi_{t}$, making numerical simulations more time-efficient than directly simulating the second-order SG equation in Eq. 2. In the next sections, we demonstrate the stability and effectiveness of this scheme for studying the long-time nonlinear dynamics of the SG equation under various perturbations.
@@ V.1.1 Single fluxon dynamics
We first apply the numerical scheme developed above to compute the dynamics of a single fluxon in an ideal long JJ, a case that has been well studied using other numerical techniques [62, 8]. Here, the dynamics of the generalized flux variable $\varphi$ are governed by the unperturbed SG equation ($\alpha=\beta=0$). The boundary conditions [3] are given by
@@
where $\eta$ is the normalized value of the $y$ component of external magentic field, and $\xi$ is the normalized value of the external current injected into the sides of the junction. In an ideal and infinitely long junction, the analytical soliton solution describing a single fluxon is given by [1]
@@
where $n$ is an integer, $x_{0}$ is the location of the fluxon at $t=0$, and $u$ is its velocity that is bounded by the Swihart velocity [9]. In Eq. 19, the Swihart velocity is normalized, $0\leq u\leq 1$. The solution in Eq. 19 corresponds to a $2\pi-$kink in $\varphi$ that travels at constant speed $u$ and is the 1+1D description of a single (2+1D) vortex trapped between the superconducting islands of the junction. One can show that the 1+1D dynamics of $\varphi$ that is given by the sine-Gordon equation can be derived from the full 2+1D electro-hydrodynamical equations that couple the superconducting order parameter and the EM field. For completeness, in Appendix A we present a rigorous derivation of this reduction.
@@
Figure 4: Comparisons on the final positions of the fluxon in the junction at the last time step ($t=T_{m}=50000$) obtained with different discretization steps $\Delta x$ and $\Delta t$. (a) shows the results from using Euler method, and (b) shows the results from DEC-QED. In all the plots from both figures, $\Delta t=0.8\Delta x$, $u=0.55$, and $L=100$.
@@
Using the expression in Eq. 19 at $t=0$ as the initial condition, we perform long-time simulations of a single fluxon in a long JJ. Figs. 3a, 3b, and 3c show the plots for $\varphi$, $\phi_{t}$, and $\phi_{x}$ respectively when there is no external bias on the boundary ($\eta\!=\!\xi\!=\!0$), implemented by a Neumann-type boundary condition for $\varphi$ at $x=0,L$. This translates to Dirichlet conditions on the edge field $\phi_{x}$. The length of the junction is $L\!=\!100$, with a total time $T_{m}\!=\!50000$, and the initial condition is given by Eq. 19 in which $u\!=\!0.55$, $x_{0}\!=\!0$, and $n\!=\!0$. The Lorentz factor, which determines the size of the soliton, is therefore $1/\sqrt{1-u^{2}}\approx 1.2$. The spacings in the space-time grid are kept at $\Delta x\!=\!0.05,$ and $\Delta t\!=\!0.04$ in order to resolve the kink. With no external bias on the boundary, the fluxon simply collides with the junction boundary, is reflected back, and becomes an anti-fluxon that travels in the opposite direction with the same velocity. Without external force or perturbations, this process repeats indefinitely - as shown in Figs. 3a-3c. Within the simulated time $(T_{m}\!=\!50000)$ the fluxon collides with the two boundaries a total of 278 times, while still maintaining its original shape and speed. By the end of the simulation, the soliton has traveled a distance that is 11583 times its size, providing an early hint for the stability of the numerical scheme even in the long-time regime. To also study the effect of boundary currents and magnetic fields, in Appendix E we present simulations for the long-time dynamics with nonzero $\eta$ and $\xi$.
@@
For a more quantitative insight into the stability of DEC-based integration, we now provide a comparative analysis to the Euler method. Although the Euler method is generally first-order accurate for linear wave equations and can be unstable at long times due to dispersion [63, 64, 65], the soliton solutions to the unperturbed sine-Gordon equation are the steady-state solutions in which the nonlinear term perfectly balances out the dispersion. Therefore, the Euler method can be used here to compare with our method. In Figs. 4a and 4b we compare the final positions of the fluxon in the JJ at the last time step $T_{m}\!=\!50000$ computed using the two methods with various discretization steps $\Delta x$ and $\Delta t$. With sufficiently small steps ($\Delta x\!=\!0.025$), the Euler method and DEC converge to the the same location for the fluxon at $t=T_{m}$ that we use as the benchmark (the blue curves in Fig. 4a and Fig. 4b) for more more coarse discretizations. As we increase $\Delta x$ (and $\Delta t$ accordingly by setting $\Delta t\!=\!0.8\Delta x$), the final location of the fluxon gradually drifts away from the benchmark, but solutions produced by DEC are noticeably closer to the benchmark than those obtained from the Euler method with the same dicretization. Overall, as $\Delta x$ increases, Euler’s method diverges from the accurate solution faster than DEC does. In Fig.4 the discrepancies are most significant with $\Delta x\!=\!0.4$, when the soliton in both methods falls behind in the number of collisions with the boundary; the soliton in the Euler method only completes $n_{c}=268$ collisions, 10 less than 278. The soliton in DEC, on the other hand, has $n_{c}=276$, only two collisions shy of the correct number.
@@ V.1.2 Fluxon-antifluxon dynamics
In general, the unperturbed sine-Gordon equation admits solutions that contain an arbitrary number of fluxons, antifluxons, and non-soliton radiation. A case that is typically investigated is the vortex-antivortex solution whose analytical form is known [1]
@@ V.2 Vortices under perturbations
Figure 6: Dynamics of a SG soliton under dissipation and external bias current. The total simulated time is $T_{m}=1000$, the initial velocity of the fluxon is $u=0.55$, and the junction length is $L=100$ with boundary coefficients $\eta=0.002$ and $\xi=0.006$. In parts (a)-(f), the initial soliton is a fluxon. The coefficients for the resistive loss and bias current (a) and (b) are [$\alpha=0.003$, $\beta=0$], (c) and (d) are [$\alpha=0.03$, $\beta=0$], and in (e) and (f) are [$\alpha=0.003$, $\beta=0.001$]. The initial soliton in parts (g) and (h) is a breather traveling in a medium with coefficients [$\alpha=0.003$, $\beta=0$].
@@ V.2.1 Solitons under resistive loss and external bias
In Figs. 6a-b, we plot the dynamics of a fluxon that has initial velocity $u\!=\!0.55$ and travels in a lossy junction with $\alpha\!=\!0.003$ and without a bias current ($\beta=0$). The fluxon slows down over time due to friction, as evidently seen from the change in the slope of the trace that the soliton makes in space-time. The fading intensity of the trace in Fig. 6a, where we plot $\phi_{t}$, also signifies the slowing down of the soliton. This is more apparent in Figs. 6c-d, where the dissipation rate is stronger with $\alpha\!=\!0.03$, forcing the soliton to quickly come to a stop without ever reaching the junction boundary. In Figs. 6e-f the dissipation rate is set to $\alpha\!=\!0.003$ as in Figs. 6a-b, but now we also include a bias current by setting $\beta\!=\!0.001$. This bias current exerts a Lorentz force on the soliton and pulls it towards the $-x$ direction if the soliton is a vortex and towards the $+x$ direction if it is an antivortex. We can see this force in action in Figs. 6e-f; the vortex is initially set to travel towards the $+x$ direction until it hits the boundary and becomes an antivortex that travels in the $-x$ direction. The antivortex slows down due to both the friction coming from the $\alpha$ term and the Lorentz force pulling it in the $+x$ direction and comes to a complete stop. It then accelerates in the opposite direction ($+x$) due to the Lorentz force until it hits the boundary and turns into a vortex again.
@@ V.2.3 Interaction between a vortex and a microshort
Mircoshorts are micro-scale (or sometimes nano-scale) short-circuits that can appear in long Josephson junctions. These short circuits happen when the insulating region becomes too narrow at some locations in the junction, causing current density to approach infinity there. They can either be unwanted defects that occur due to fabrication imperfections or intentional design decisions [66]. Therefore, precise modeling of vortex interactions near microshorts is important. The modified sine-Gordon equation that takes into account the presence of microshorts is given by
@@
We therefore see that for $\varphi=2\pi n$ with $n$ being an arbitrary integer, Eq. 23 vanishes and there is no energy stored in the short. If a fluxon is in the vicinity of the microshort, $\varphi$ deviates from the bulk value and now $H_{s}>0$. The short draws energy directly from the vortex and causes it to slow down, or equivalently the vortex experiences a repulsion as it approaches the short. The dynamics of this process is shown in Fig. 8. We consider a junction whose length is $L=40$, has a resistive loss rate of $\alpha=0.005$ and is subjected to a bias current $\beta$. A fluxon is initially prepared at terminal velocity $u_{t}$ (velocity at which friction and Lorentz force cancel each other) at $x_{0}=10$, sufficiently far away from and moving towards a microshort at $x_{s}=-10$. In Fig. 8b, where we set $u_{t}=0.3$ (which corresponds to $\beta=0.002$), the fluxon is repelled by the short. After the first repulsion, the fluxon travels in the opposite direction until it is slowed down to a complete stop by both resistance and the Lorentz force. The Lorentz force, which always pulls the fluxon in the $-x$ direction, then accelerates the fluxon towards the short again. As seen in Fig. 8b, the fluxon bounces off the mircoshort multiple times, losing energy in the process due to resistance, and eventually oscillates inwards into the pinning position, which is an equilibrium point $x_{e}$ where the Lorentz force and the repulsion from the short balance out each other (see inset of Fig. 8b). Note that when the fluxon is repelled by the microshort, it does not turn into an antifluxon, unlike what happens when it collides with the boundary.
If the initial kinetic energy of the vortex is sufficiently large, it can escape pinning [1]. This scenario is shown in Fig. 8c, in which the initial speed of the soliton is $u_{t}=0.618$ and corresponds to a bias current $\beta=0.005$. The fluxon simply passes through the microshort with negligible changes in kinetic energy.
Microshorts are the extreme examples of when Josephson junctions have varying critical current across their length. More generally, the thickness of the insulating region separating the superconducting islands of a long junction may be nonuniform, which may affect the structure and the dynamics of fluxons as they travel along the junction. In Appendix F we investigate such a situation, where we consider a fluxon traveling in junctions that have one or a few constricted regions that force the fluxon to slow down by emitting radiation. In particular, we observe signatures of Cherenkov radiation [67, 45] when a fluxon arrives at a constriction with an incident velocity larger than the Swihart velocity in that constricted region.

### v3

@@ IV Discretization of 1+1D field theory with DEC-QED
In this section, we discuss a discretization procedure for the 1+1D field equations to enable their symplectic integration. Although in this section we will focus on the derivation of the coarse-grained version of the perturbed sine-Gordon equation in space-time, the process for coarse-graining other nonlinear field equations follows a similar procedure (see e.g. Ref. [dec-qed]). We consider a primal rectangular grid $M$ that discretizes the space-time computational domain, in which $\Delta x$ and $\Delta t$ are spacings between the neighboring grid points along $x$ and $t$, respectively. We label the vertices, edges, and elemental faces of this mesh as $v$, $e$ and $f$, respectively. Additionally, consider a dual mesh whose vertices are the circumcenters of the rectangular cells $f$ of the primal grid. Each edge of this dual mesh is constructed by connecting the two dual nodes corresponding to two neighboring primal faces (faces that share a primal edge). The dual faces, edges, and nodes are labeled $v^{\dagger}$, $e^{\dagger}$, and $f^{\dagger}$, respectively. Fig. 2a presents a schematic of the spacetime grid, where the primal mesh is colored orange and the dual mesh is in green.
By introducing the 1+1D Minkowski metric tensor $g_{\mu\nu}$ [lee2006riemannian, dirac1996GR], Eq. 3 can be conveniently rewritten as
@@
$0$ $=\int_{v^{\dagger}}(g^{\mu\nu}\partial_{\mu}\partial_{\nu}\varphi+\alpha\partial_{t}\varphi+\sin\varphi+\beta)dA$ (13)
$=\!\int_{\partial v^{\dagger}}\!\!\hat{n}_{\mu}\!\cdot\!\left(\nabla\varphi\right)^{\mu}ds+\left(\alpha\partial_{t}\varphi\!+\!\sin\varphi\!+\!\beta\right)\!\Delta A$
@@
where $\bar{\phi_{t}}$ is the average of the field values on the two $\phi_{t}$ edges connected to $v$. Given initial conditions on either $\phi_{t}$ and $\varphi$ everywhere in space or $\phi_{t}$ and $\phi_{x}$ everywhere combined with the value for $\varphi$ at the boundary, Eq. 14 can be used to propagate the system in time to study the evolution of $\phi_{t}$, $\phi_{x}$. The procedure is as follows: by labeling each primal vertex using their discrete spatial and temporal indices $i$ and $j$ and each primal edge by the two vertices connected to it, Eq. 14 when applied to a vertex $v(i,j)$ can be rewritten as
@@
$\hskip-8.67204pt\varphi(i,j\!+\!1)=\phi_{t}[(i,j),(i,j\!+\!1)]\!+\!\varphi(i,j),$ (16)
$\hskip-8.67204pt\phi_{x}[(i,j\!+\!1),\!(i\!+\!1,j\!+\!1)]=\varphi(i\!+\!1,j\!+\!1)\!-\!\varphi(i,j\!+\!1).$ (17)
Eqs. IV, 16, and 17, along with initial and boundary conditions allow us to determine the fields $\phi_{t},\phi_{x},$ and $\varphi$ everywhere on the spacetime grid. Note that Eq. 14 is effectively a first-order equation in terms of the edge fields $\phi_{x}$ and $\phi_{t}$, making numerical simulations more time-efficient than directly simulating the second-order SG equation in Eq. 3. In the next sections, we demonstrate the stability and effectiveness of this scheme for studying the long-time nonlinear dynamics of the SG equation under various perturbations.
@@ V.1.1 Single fluxon dynamics
We first apply the numerical scheme developed above to compute the dynamics of a single fluxon in an ideal long JJ, a case that has been well studied using other numerical techniques [nakajima1974, krasnov2020josephson]. Here, the dynamics of the generalized flux variable $\varphi$ are governed by the unperturbed SG equation ($\alpha=\beta=0$). The boundary conditions [ustinov1998solitons], which are typically applied to inline junctions, are given by
@@
where $\eta$ is the normalized value of the $y$ component of external magnetic field, and $\xi$ is the normalized value of the external current injected into the sides of the junction. In an ideal and infinitely long junction, the analytical soliton solution describing a single fluxon is given by [mclaughlin1978_SG]
@@
where $n$ is an integer, $x_{0}$ is the location of the fluxon at $t=0$, and $u$ is its velocity that is normalized by the Swihart velocity [swihart1961]. In Eq. 20, the Swihart velocity is normalized, $0\leq u\leq 1$. The solution in Eq. 20 corresponds to a $2\pi$-kink in $\varphi$ that travels at constant speed $u$ and is the 1+1D description of a single (2+1D) vortex trapped between the superconducting islands of the junction. One can show that the 1+1D dynamics of $\varphi$ that is given by the sine-Gordon equation can be derived from the full 2+1D electro-hydrodynamical equations that couple the superconducting order parameter and the EM field. For completeness, in Appendix A we present a rigorous derivation of this reduction.
@@
Figure 4: Comparisons on the final positions of the fluxon in the junction at the last time step ($t=T_{m}=50000$) obtained with different discretization steps $\Delta x$ and $\Delta t$. (a) shows the results from using Euler method, and (b) shows the results from DEC-QED. In all the plots from both figures, $\Delta t=0.8\Delta x$, $u=0.55$, and $L=100$. The inset in (b) shows the difference $\Delta D$ in the distance traveled by the soliton calculated with different values of $\Delta x$ to the distance $D$ covered by the soliton simulated with $\Delta x=0.025$.
@@
Using the expression in Eq. 20 at $t=0$ as the initial condition, we perform long-time simulations of a single fluxon in a long JJ. Figs. 3a, 3b, and 3c show the plots for $\varphi$, $\phi_{t}$, and $\phi_{x}$ respectively when there is no external bias on the boundary ($\eta\!=\!\xi\!=\!0$), implemented by a Neumann-type boundary condition for $\varphi$ at $x=0,L$. This translates to Dirichlet conditions on the edge field $\phi_{x}$. The length of the junction is L=100L\!=\!100, with a total time $T_{m}\!=\!50000$, and the initial condition is given by Eq. 20 in which u=0.55u\!=\!0.55, $x_{0}\!=\!0$, and n=0n\!=\!0. The Lorentz factor, which determines the size of the soliton, is therefore $1/\sqrt{1-u^{2}}\approx 1.2$. The spacings in the space-time grid are kept at $\Delta x\!=\!0.05,$ and $\Delta t\!=\!0.04$ in order to resolve the kink. With no external bias on the boundary, the fluxon simply collides with the junction boundary, is reflected back, and becomes an anti-fluxon that travels in the opposite direction with the same velocity. Without external force or perturbations, this process repeats indefinitely - as shown in Figs. 3a-3c. Within the simulated time $(T_{m}\!=\!50000)$ the fluxon collides with the two boundaries a total of 278 times, while still maintaining its original shape and speed. By the end of the simulation, the soliton has traveled a distance that is 11583 times its size, providing an early hint for the stability of the numerical scheme even in the long-time regime. To also study the effect of boundary currents and magnetic fields, in Appendix E we present simulations for the long-time dynamics with nonzero $\eta$ and $\xi$.
@@
For a more quantitative insight into the stability of DEC-based integration, we now provide a comparative analysis to the Euler method. Although the Euler method is generally first-order accurate for linear wave equations and can be unstable at long times due to dispersion [shlager1993relative, petropoulos1994phase, hyatt2025multiple], the soliton solutions to the unperturbed sine-Gordon equation are the steady-state solutions in which the nonlinear term perfectly balances out the dispersion. Therefore, the Euler method can be used here to compare with our method. In Figs. 4a and 4b we compare the final positions of the fluxon in the JJ at the last time step $T_{m}\!=\!50000$ computed using the two methods with various discretization steps $\Delta x$ and $\Delta t$. With sufficiently small steps ($\Delta x\!=\!0.025$), the Euler method and DEC converge to the the same location for the fluxon at $t=T_{m}$ that we use as the benchmark (the blue curves in Fig. 4a and Fig. 4b) for more more coarse discretizations. As we increase $\Delta x$ (and $\Delta t$ accordingly by setting $\Delta t\!=\!0.8\Delta x$), the final location of the fluxon gradually drifts away from the benchmark, but solutions produced by DEC are noticeably closer to the benchmark than those obtained from the Euler method with the same dicretization. Overall, as $\Delta x$ increases, Euler’s method diverges from the accurate solution faster than DEC does. In Fig. 4 the discrepancies are most significant with $\Delta x\!=\!0.4$, when the soliton in both methods falls behind in the number of collisions with the boundary; the soliton in the Euler method only completes $n_{c}=268$ collisions, 10 less than 278. The soliton in DEC, on the other hand, has $n_{c}=276$, only two collisions shy of the correct number. In the inset in Fig. 4b we show as a function of $1/\Delta x$ the difference $\Delta D$ in the distance traveled by the soliton with respect to the distance $D$ that the soliton traveled when a sufficiently small $\Delta x=0.025$ was used; even for $\Delta x=0.4$ – a discretization step 16 times larger than the benchmark step size – the relative error in the final position of the soliton is only $\Delta D/D\approx 3\times 10^{-3}$ after having traveled a distance that is 11583 times its size.
@@ V.1.2 Fluxon-antifluxon dynamics
In general, the unperturbed sine-Gordon equation admits solutions that contain an arbitrary number of fluxons, antifluxons, and non-soliton radiation. A case that is typically investigated is the vortex-antivortex solution whose analytical form is known [mclaughlin1978_SG]
@@ V.2 Vortices under perturbations
Figure 6: Dynamics of a SG soliton under dissipation and external bias current. The total simulated time is $T_{m}=1000$, the initial velocity of the fluxon is $u=0.55$, and the junction length is $L=100$ with boundary coefficients $\eta=0.002$ and $\xi=0.006$. In parts (a)-(f), the initial soliton is a fluxon. The coefficients for the resistive loss and bias current (a) and (b) are $\alpha=0.003$, $\beta=0$], (c) and (d) are $\alpha=0.03$, $\beta=0$], and in (e) and (f) are $\alpha=0.003$, $\beta=0.001$]. The initial soliton in parts (g) and (h) is a breather traveling in a medium with coefficients $\alpha=0.003$, $\beta=0$].
@@ V.2.1 Solitons under resistive loss and external bias
In Figs. 6a-b, we plot the dynamics of a fluxon that has initial velocity u=0.55u\!=\!0.55 and travels in a lossy junction with $\alpha\!=\!0.003$ and without a bias current ($\beta=0$). The fluxon slows down over time due to friction, as evidently seen from the change in the slope of the trace that the soliton makes in space-time. The fading intensity of the trace in Fig. 6a, where we plot $\phi_{t}$, also signifies the slowing down of the soliton. This is more apparent in Figs. 6c-d, where the dissipation rate is stronger with $\alpha\!=\!0.03$, forcing the soliton to quickly come to a stop without ever reaching the junction boundary. In Figs. 6e-f the dissipation rate is set to $\alpha\!=\!0.003$ as in Figs. 6a-b, but now we also include a bias current by setting $\beta\!=\!0.001$. This bias current exerts a Lorentz force on the soliton and pulls it towards the x-x direction if the soliton is a vortex and towards the +x+x direction if it is an antivortex. We can see this force in action in Figs. 6e-f; the vortex is initially set to travel towards the +x+x direction until it hits the boundary and becomes an antivortex that travels in the x-x direction. The antivortex slows down due to both the friction coming from the $\alpha$ term and the Lorentz force pulling it in the +x+x direction and comes to a complete stop. It then accelerates in the opposite direction (+x+x) due to the Lorentz force until it hits the boundary and turns into a vortex again.
@@ V.2.3 Interaction between a vortex and a microshort
Mircoshorts are micro-scale (or sometimes nano-scale) short-circuits that can appear in long Josephson junctions. These short circuits happen when the insulating region becomes too narrow at some locations in the junction, causing current density to approach infinity there. They can either be unwanted defects that occur due to fabrication imperfections or intentional design decisions [price2010vortexqbit]. Therefore, precise modeling of vortex interactions near microshorts is important. The modified sine-Gordon equation that takes into account the presence of microshorts is given by
@@
We therefore see that for $\varphi=2\pi n$ with $n$ being an arbitrary integer, Eq. 24 vanishes and there is no energy stored in the short. If a fluxon is in the vicinity of the microshort, $\varphi$ deviates from the bulk value and now $H_{s}>0$. The short draws energy directly from the vortex and causes it to slow down, or equivalently the vortex experiences a repulsion as it approaches the short. The dynamics of this process is shown in Fig. 8. We consider a junction whose length is $L=40$, has a resistive loss rate of $\alpha=0.005$ and is subjected to a bias current $\beta$. A fluxon is initially prepared at terminal velocity $u_{t}$ (velocity at which friction and Lorentz force cancel each other) at $x_{0}=10$, sufficiently far away from and moving towards a microshort at $x_{s}=-10$. In Fig. 8b, where we set $u_{t}=0.3$ (which corresponds to $\beta=0.002$), the fluxon is repelled by the short. After the first repulsion, the fluxon travels in the opposite direction until it is slowed down to a complete stop by both resistance and the Lorentz force. The Lorentz force, which always pulls the fluxon in the x-x direction, then accelerates the fluxon towards the short again. As seen in Fig. 8b, the fluxon bounces off the mircoshort multiple times, losing energy in the process due to resistance, and eventually oscillates inwards into the pinning position, which is an equilibrium point $x_{e}$ where the Lorentz force and the repulsion from the short balance out each other (see inset of Fig. 8b). Note that when the fluxon is repelled by the microshort, it does not turn into an antifluxon, unlike what happens when it collides with the boundary.
If the initial kinetic energy of the vortex is sufficiently large, it can escape pinning [mclaughlin1978_SG]. This scenario is shown in Fig. 8c, in which the initial speed of the soliton is $u_{t}=0.618$ and corresponds to a bias current $\beta=0.005$. The fluxon simply passes through the microshort with negligible changes in kinetic energy.
Microshorts are the extreme examples of when Josephson junctions have varying critical current across their length. More generally, the thickness of the insulating region separating the superconducting islands of a long junction may be nonuniform, which may affect the structure and the dynamics of fluxons as they travel along the junction. In Appendix F we investigate such a situation, where we consider a fluxon traveling in junctions that have one or a few constricted regions that force the fluxon to slow down by emitting radiation. In particular, we observe signatures of Cherenkov radiation [akhmediev1995cherenkov, goldobin1998cherenkov] when a fluxon arrives at a constriction with an incident velocity larger than the Swihart velocity in that constricted region.